# A physically modelled selection function for compact binary mergers in the LIGO-Virgo O3 run and beyond

### v1

@@ 1 Introduction
Gravitational wave (GW) astronomy has made a number of breakthroughs in recent years, with the first detection of merging black holes by the interferometric gravitational-wave detector network on 14 September 2015 [1]. Coalescences of compact binaries (CBC), such as black hole or neutron star binaries, are a particularly interesting source for broad-band detectors, since the signal from an inspiral is quite strong compared to other types of GWs, and the binary system properties can be measured with high precision for some parameters of astrophysical interest, due to the large number of signal cycles observable in the sensitive frequency band [2]. Many unanswered questions about such sources remain, despite recent LIGO-Virgo observations [3], such as the origin of GW-emitting black hole binary (BBH) systems, which have properties very different from black holes detected via X-ray emission [4] (though compare [5]).
@@
Many formation channels and mechanisms have been proposed for the BBH observed via GW emission (see [6] for an overview). Among the most intensively studied are isolated binary evolution (e.g. [7, 8], and dynamical binary formation in globular or other stellar clusters (e.g. [9, 10]) or around galactic centres (e.g. [11]). Each of these channels may have distinctive signatures in the distribution of merging binaries over masses, spins and redshift, though subject to many theoretical and modelling uncertainties (e.g. [12]; see [13, 14, 15] for more detailed discussion). Currently, no single formation channel appears capable of explaining the entire population [16], implying that a more complex mixture of channels will eventually be crucial to properly describe observations. The current detector network is only sensitive over a limited range of redshift, thus any population evolution over cosmic history is so far not unambiguously observed [17, 18, 19, 20, 21]. Since various formation channels are expected to have different dependences on redshift [22, 23, 24], tracking such evolution will become increasingly important to untangle them as detector sensitivity increases (e.g. [25]).
@@
The set of binary mergers that we observe comes with a significant bias as compared to the actual astrophysical BBH population, because of selection effects, which are crucial to take into account in population studies. The most direct selection effect, and the one currently limiting our ability to probe cosmic evolution, is that nearby (low redshift) sources are more likely to be detected than far (high redshift) ones. Nearly as important is the dependence of signal amplitude on binary masses: for the inspiral phase of the signal, this varies as (a power of) the chirp mass $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$, while for the later merger-ringdown phase the binary total mass has the most influence. The spins of binary components also affect the GW signal detectability[26, 27], primarily via a longer or shorter signal duration with the merger at higher (lower) frequency for spins aligned (anti-aligned) with the orbit. Spins in the plane of the orbit, which give rise to precession effects, also affect detection, primarily since the existing searches use templates with only orbit-aligned component spins, which therefore cannot fully match signals with precession (e.g. [28]); similar considerations apply to non-dominant multipole (“higher mode”) emission which is not included in current template banks [29, 30]. Signals from binaries with non-negligible residual orbital eccentricity likewise will be less efficiently recovered [31, 32].
@@
The interplay of search sensitivity with these astrophysical properties of binaries makes it crucial to have an accurate quantitative estimate of selection effects when carrying out population analysis. This is provided by an estimate of the probability of detection $P_{\mathrm{det}}$, or selection function, of binary mergers: an accurate estimate enables measurement of the rates of binary merger events, and reconstruction of their population distribution as a function of component masses, spins and redshift [13, 14]. Events are considered to be detected on the basis of their statistical significance assigned by search pipelines, i.e. matched-filtering algorithms looking for signals in the data [33, 34, 35, 36, 37]. The significance is quantified by the false alarm rate (FAR), which is the expected rate of candidates in the pipeline resulting from detector noise which are ranked equal to or higher than the candidate under consideration. A set of candidates of well defined purity, in terms of the expected number of noise events, may be obtained by imposing a pre-defined maximum threshold on FAR. BBH candidates have typically been considered for population analysis if their FAR is below a threshold $1/\mathrm{yr}$; in the case of several pipelines, the most significant result for a given candidate has to be under the threshold [13, 14].
@@
where $P_{\mathrm{det}}(\theta)$ stands for the probability that an event with parameters $\theta$ would be detected (found) by a particular search or a combination of searches. For binary population analyses, $P_{\mathrm{det}}$ is estimated by simulating CBC signals from a reference BBH population, adding them to real detector data and analysing the “injected” data with search pipelines to record which ones are detected [14]. Since the previous integral cannot be evaluated analytically, importance sampling is used instead to estimate it using a weighted Monte Carlo integral over found injections, i.e. those satisfying the chosen search significance threshold:
@@
in which $N_{\mathrm{inj}}$ is the total number of injections, $N_{\mathrm{det}}$ are the found (detected) injections, and $p_{\mathrm{draw}}$ is a fiducial distribution from which the samples are drawn [40, 39].
Importance sampling is a convenient approach since it only requires one pre-determined injection analysis campaign to be performed by searches, which is then used to estimate selection effects for any population models considered. However, this procedure comes with uncertainty in the estimation of the selection integral: the precision of the estimation depends on how well the proposal model $\pi$ matches the form of $p_{\mathrm{draw}}$. There must be enough samples drawn and detected by search pipelines so that this uncertainty does not alter the shape of the population model posterior [39, 41]. The precision may be quantified via the effective number of independent draws (sample size) ($N_{\mathrm{eff}}$) that contribute to estimate $\xi$ for a given model. This effective sample size is defined as the square of the estimated fraction $\hat{\xi}$ divided by its variance due to Monte Carlo uncertainty: in terms of the weights $w_{j}=\pi(\theta_{j})/p_{\mathrm{draw}}(\theta_{j})$, $N_{\mathrm{eff}}$ varies as
@@
Then depending on the model $\pi(\theta|\Lambda)$, for a fixed injection set $N_{\mathrm{eff}}$ may or may not be high enough to obtain an accurate population inference; as a general guide $N_{\mathrm{eff}}>4N_{\mathrm{det}}$ has been required [39, 41]. Then, if we discount the possibility of requiring more injections to be analyzed ad-hoc, in general at high computational cost, a large class of models cannot be studied using a generic (broadly/uniformly distributed) injection set. For example, population models where $\pi(\theta|\Lambda)$ is strongly peaked around a specific point while the density of injections $p_{\mathrm{draw}}(\theta)$ is not (e.g. [42]); or in general, where some $w_{i}$ are $\gg 1$ (see also [43]).
As an alternative, using an analytical function – or in general, a continuous function – that accurately describes the probability of detection of BBH signals will make it possible to estimate $P_{\mathrm{det}}$ without such restrictions on the model and original injection sampling distributions. Some efforts were already made towards describing selection effects in this way; to begin with, the dependence on masses and spins of the selection function can be modelled via “semi-analytic” methods with corrections applied by various methods. The interferometer network sensitivity is approximated by evaluating the expected (optimal) signal-to-noise ratios (SNRs) of signals for fixed, ideal detector power spectral densities (PSDs). The probability of detection is then obtained by assuming that the network is sensitive to CBCs with SNR above a fiducial threshold value [44, 45], which may be tuned to approximate the number of actual signals, or injections detected by search pipelines [46]. However, a simple optimal SNR threshold is not necessarily an accurate approximation for $P_{\mathrm{det}}$, due to many factors which violate the assumptions behind the “semi-analytic” calculation. Beyond the effect of the noise realization on matched filter SNR, for instance, the number of detectors observing and their sensitivity varies over time; also, detector noise is not perfectly described by the PSDs, since that would imply it to be perfectly Gaussian and stationary. In reality there are excess noise transients (“glitches”, see e.g. [47, 48]) that may be partly suppressed by search pipelines employing additional quantities beyond the SNR, and other non-ideal noise artefacts. In general the searches might have different ranking statistics that have additional dependences beyond the SNR [49, 50, 51], for example explicit dependence on masses and spins.
@@
The strategy we pursue here to overcome this limitation is to fit a continuous and physically motivated $P_{\mathrm{det}}$ function over distance, masses and even spins. While the basic functional form is derived from the known theoretical dependence of SNR on distance and chirp mass, we introduce corrections to account for non-ideal detector behaviour, and for higher order effects of masses and spins on an event’s detectability in actual search pipelines. The fit is adjusted to maximize the likelihood of found injections in the O3 sensitivity data release corresponding to GWTC-3 [54, 3].
@@ 2 Methods
LIGO-Virgo O3 BBH search sensitivity estimates [54] were used for the analysis. This data release contains sets of simulated signals (injections) from binary black holes that can be used to estimate the search sensitivity throughout the GWTC-3 catalog from the LVK (LIGO-Virgo-KAGRA) Collaboration [55].
@@ 2.1 Sigmoid distance dependence model
Considering a source of fixed intrinsic parameters (binary component masses and spins), the dependence on distance of search sensitivity is mainly determined by the effects of angular parameters giving the sky position and orientation of the source relative to a detector network (assumed to have stationary sensitivity). If the angular parameters were also fixed, then the amplitude and expected SNR are simply inversely proportional to distance: then, assuming the source is detected if SNR is above a given threshold, $P_{\mathrm{det}}(d_{L})$ would be almost 1 for some range of distances up to a given point, and then close to 0 for larger distances, as represented in Fig. 1a. The detection probability has the shape of a sigmoid instead of a step function because of noise in the detector: in an actual search, detection will be determined by the matched filter SNR, a function of the data which contains a stochastic term resulting from the detector noise realization in addition to the expected SNR.
@@
Since the sources’ positions and orientations are a priori unknown, and isotropically distributed throughout the sky, we need to consider all the possible values of the angular parameters and average over them. As a result of this marginalization, the functional form of the search sensitivity takes a form more similar to Fig. 1b. These effects on search sensitivity were quantified by Finn & Chernoff (FC [56]), predicting the form of the probability of detection function over distance by considering geometrical factors in the expected SNR for the case of only one detector with fixed sensitivity [56].
Although the true functional dependence of $P_{\mathrm{det}}(d_{L})$ in a realistic case will have many corrections relative to this idealized calculation, our fundamental model assumption is that such corrections are small: thus, to proceed, we use as a starting point a parameterized sigmoid function that is able to fit the FC data (Table I, [56]) faithfully, while also allowing for minor deviations. The shape of the sigmoid is controlled by four parameters $\varepsilon_{\max},\alpha,\delta$ and $\gamma$; out of which $\alpha=2.05$ is fixed to the value obtained in a direct fit to FC data and the other 3 will be free model parameters. The probability of detection, for a binary with given intrinsic parameters, thus reads as
@@
Having modelled the dependence on amplitude and distance for fixed intrinsic binary parameters, we then add a dependence on these intrinsic parameters by writing $d_{\mathrm{mid}}(m_{1z},m_{2z})$ and $\varepsilon_{\max}(m_{1z},m_{2z})$ as a function of the redshifted masses (we will also allow for component spin dependence in Sec. 4). In the low mass limit, $d_{\mathrm{mid}}$ should look somewhat like $d_{\mathrm{mid}}\sim D_{0}f_{d}\mathcal{M}_{z}^{5/6}$, where $D_{0}$ would be sort of the average sensitive distance over the whole run and detectors; i.e.  the sensitive distance should scale with $\mathcal{M}_{z}^{5/6}$ for the simplest inspiral model, where the redshifted chirp mass is given by $\mathcal{M}_{z}=\eta^{3/5}M_{z}$; $M_{z}=m_{1z}+m_{2z}$ stands for the total redshifted mass of the binary system and $\eta=m_{1z}m_{2z}/M_{z}^{2}$ the symmetric mass ratio. The remaining dependence on masses, corresponding to corrections to this lowest order inspiral SNR power law behaviour, is then to be fitted by a function $f_{d}(M_{z},\eta)$. Taking into account that $\varepsilon_{\max}$ should not be higher than 1 and always positive, we take a functional form $\varepsilon_{\max}\sim 1-\exp(f_{\varepsilon})$, with $f_{\varepsilon}$ a function of intrinsic parameters.
At this point, the functions providing mass dependence to $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ could be any arbitrary combination of $M_{z}$ and $\eta$. In order to gain more insight into the needed corrections, mass binned fits to the detected injections from the GWTC-3 catalog were performed.
@@ 2.2 Mass binned fits
In order to gain insight into the functional forms of $f_{d}$ and $f_{\varepsilon}$ that will be able to fit injection results over the BBH mass space, we first fit our analytical model of $P_{\mathrm{det}}$ as a function of distance for ‘fixed’ masses or, rather, in mass bins, by numerically maximizing the likelihood of the injection campaign search results. [^1] Code to evaluate our $P_{\mathrm{det}}$ model and fit its parameters to injection results, and the resulting maximum likelihood parameter values, are publicly available at https://github.com/AnaLorenzoMedina/cbc_pdet. In this binned fit procedure, rather than being continuous functions, $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ are treated as scalar parameters, fitted separately for each mass bin. For a given mass bin labelled by $M$ we use a inhomogeneous Poisson likelihood for the found injections (the ones with FAR $\leq 1$/yr, similarly to [14]):
@@
$\mu^{M}\equiv N_{\text{exp}}=\int_{0}^{\infty}\lambda^{M}(d_{L})\,\text{d}d_{L}\,.$ (7)
Although $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ vary between bins, we use the same shape parameter values $\delta$ and $\gamma$ over the whole mass range, since we assume $P_{\mathrm{det}}$ should have the same shape regardless of the mass. We then perform a fit in two alternating stages. In the ‘sigmoid fitting’ stage, $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ are held fixed for all bins and the sigmoid shape is fitted by maximizing the total log likelihood $\sum_{M}\ln\mathcal{L}^{M}$ over $\delta$ and $\gamma$. In the ‘mass fitting’ stage, $\delta$ and $\gamma$ are instead fixed and the values of $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ are fitted separately for each bin by maximizing $\ln\mathcal{L}^{M}$. The two stages are alternated until the total likelihood has converged to a global maximum. (For the actual optimization at each stage, the Nelder-Mead simplex algorithm was used, which is a typical direct search method for multidimensional unconstrained minimization [58].)
Figure 2 shows the resulting optimized values of $d_{\mathrm{mid}}$, having scaled out a factor of $\mathcal{M}_{z}^{5/6}$. While all the values are comparable in magnitude, we find strong linear and quadratic dependences on $M_{z}$ at higher $\eta$, decreasing as the symmetric mass ratio gets smaller. Taking into account this behaviour, we first proposed the following expression to describe the global mass dependence:
$d_{\mathrm{mid}}(m_{1z},m_{2z})=D_{0}\mathcal{M}_{z}^{5/6}\left[1+a_{10}M_{z}+a_{20}M_{z}^{2}+(a_{01}+a_{11}M_{z}+a_{21}M_{z}^{2})(1-4\eta)\right],$ (8)
in which we are separating out the dependence away from equal masses ($\eta=1/4$) where there may be higher uncertainty, treating this as a correction for binaries with unequal masses. This first fit was already used in a population analysis to investigate the evolution of the BBH population over redshift [21]. However, this formula for $d_{\mathrm{mid}}$ has the disadvantage that it may become negative, which would be physically disallowed, so our current :
@@
Figure 2: $d_{\mathrm{mid}}$ optimized values from the mass binned fits.
Figure 3: $\varepsilon_{\max}$ optimized values from the mass binned fits: the color bar shows the number of detected events (injections) in each mass bin.
The optimized values of $\varepsilon_{\max}$ (detection probability in the limit of zero distance) are displayed in Fig. 3. Here, we allow $\varepsilon_{\max}$ to be greater than 1 in order to allow for small number statistic fluctuations affecting the number of of injections performed within a mass bin; we see that such (in principle unphysical) $\varepsilon_{\max}$ values are mainly confined to bins with small counts of found injections. The dependence looks simplest in linear total mass, therefore to make sure the specifications stated previously are met, we chose the following function:

### v2

@@ 1 Introduction
Gravitational wave (GW) astronomy has made a number of breakthroughs in recent years, with the first detection of merging black holes by the interferometric gravitational-wave detector network on 14 September 2015 [1]. Coalescences of compact binaries (CBC), such as black hole or neutron star binaries, are a particularly interesting source for broad-band detectors, since the signal from an inspiral is quite strong compared to other types of GWs, and the binary system properties can be measured with high precision for some parameters of astrophysical interest, due to the large number of signal cycles observable in the sensitive frequency band [2]. With the current catalog of close to 100 observed binaries [3] many unanswered questions about such sources remain, such as the origin of GW-emitting black hole binary (BBH) systems, it being debatable whether they have properties different from [4] or similar to [5] black holes detected via X-ray emission.
@@
Many formation channels and mechanisms have been proposed for the BBH observed via GW emission (see [6] for an overview). Among the most intensively studied are isolated binary evolution (e.g. [7, 8], and dynamical binary formation in globular or other stellar clusters (e.g. [9, 10]) or around galactic centres (e.g. [11]). Each of these channels may have distinctive signatures in the distribution of merging binaries over masses, spins and redshift, though subject to many theoretical and modelling uncertainties (e.g. [12]; see [13, 14, 15] for more detailed discussion). Currently, no single formation channel appears capable of explaining the entire population [16], implying that a more complex mixture of channels will eventually be crucial to properly describe observations. The current detector network is only sensitive over a limited range of redshift: thus, although some studies find hints of population evolution over cosmic history, such findings are so far strongly dependent on the choice of methods [17, 18, 19, 20, 21]. Since various formation channels are expected to have different dependencies on redshift [22, 23, 24], tracking such evolution will become increasingly important to untangle them as detector sensitivity increases (e.g. [25]).
@@
The set of binary mergers that we observe comes with a significant bias as compared to the actual astrophysical BBH population, because of selection effects, which are crucial to take into account in population studies (see e.g. [26] for an early review). The most direct selection effect, and the one currently limiting our ability to probe cosmic evolution, is that nearby (low redshift) sources are more likely to be detected than far (high redshift) ones. Nearly as important is the dependence of signal amplitude on binary masses: for the inspiral phase of the signal, this varies as (a power of) the chirp mass $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$, while for the later merger-ringdown phase the binary total mass has the most influence. The spins of binary components also affect the GW signal detectability [27, 28], primarily via a longer or shorter signal duration with the merger at higher (lower) frequency for spins aligned (anti-aligned) with the orbit. Spins in the plane of the orbit, which give rise to precession effects, also affect detection, primarily since the existing searches use templates with only orbit-aligned component spins, which therefore cannot fully match signals with precession (e.g. [29]); similar considerations apply to non-dominant multipole (“higher mode”) emission which is not included in current template banks [30, 31]. Signals from binaries with non-negligible residual orbital eccentricity will also be less efficiently recovered [32, 33]; studies of eccentricity in combination with other effects will eventually be required, with accurate waveform modelling as a technical limiting factor (see e.g. discussions in [34]).
@@
The interplay of search sensitivity with these astrophysical properties of binaries makes it crucial to have an accurate quantitative estimate of selection effects when carrying out population analysis. This is provided by an estimate of the probability of detection $P_{\mathrm{det}}$, or selection function, of binary mergers: an accurate estimate enables measurement of the rates of binary merger events, and reconstruction of their population distribution as a function of component masses, spins and redshift [13, 14].
Events are considered to be detected on the basis of their statistical significance assigned by search pipelines, i.e. matched-filtering algorithms looking for signals in the data [35, 36, 37, 38, 39]. Each candidate event has a ranking statistic value calculated as a predefined function of the event parameters. The significance may be quantified by the false alarm rate (FAR), which is the expected rate of candidates in the pipeline resulting from detector noise which are ranked equal to or higher than the candidate under consideration. A set of candidates of well defined purity, in terms of the expected number of noise events, may be obtained by imposing a pre-defined maximum threshold on FAR. BBH candidates have typically been considered for population analysis if their FAR is below a threshold $1/\mathrm{yr}$; in the case of several pipelines, one may apply the threshold to the most significant result between pipelines for a given candidate [13, 14]. [^1] Other methods of quantifying event significance by pooling multiple pipeline results may be considered [40, 41].
@@
where $P_{\mathrm{det}}(\theta)$ stands for the probability that an event with parameters $\theta$ would be detected (found) by a particular search or a combination of searches. For binary population analyses, $P_{\mathrm{det}}$ is estimated indirectly by simulating CBC signals from a reference BBH population, adding them to real detector data and analysing the “injected” data with search pipelines to record which ones are detected [14]. For the detection fraction to be well defined, we must restrict the time of arrival (considered as an event parameter) to periods where any detectors were observing, and effectively calculate average detectability over such periods.
Since the previous integral cannot be evaluated analytically, importance sampling may be used instead to estimate it using a weighted Monte Carlo integral (see e.g. [44]) over found injections, i.e. those satisfying the chosen search significance threshold:
@@
in which $N_{\mathrm{inj}}$ is the total number of injections, $j=1,\ldots,N_{\mathrm{det}}$ labels the found (detected) injections, and $p_{\mathrm{draw}}$ is a fiducial distribution from which the samples are drawn [45, 43].
Importance sampling is a convenient approach since it only requires one pre-determined injection analysis campaign to be performed by searches, which is then used to estimate selection effects for any population models considered. However, this procedure comes with uncertainty in the estimation of the selection integral: the precision of the estimation depends on how well the proposal model $\pi$ matches the form of $p_{\mathrm{draw}}$. There must be enough samples drawn and detected by search pipelines so that this uncertainty does not alter the shape of the population model posterior [43, 46, 47]. The precision may be quantified via the effective number of independent draws (sample size), $N_{\mathrm{eff}}$, that contribute to estimate $\xi$ for a given model. This effective sample size is defined as the square of the estimated fraction $\hat{\xi}$ divided by its variance due to Monte Carlo uncertainty [47]; in terms of the weights $w_{j}=\pi(\theta_{j})/p_{\mathrm{draw}}(\theta_{j})$, $N_{\mathrm{eff}}$ varies as
@@
Then depending on the model $\pi(\theta|\Lambda)$, for a fixed injection set $N_{\mathrm{eff}}$ may or may not be high enough to obtain an accurate population inference; as a general guide $N_{\mathrm{eff}}>4N_{\mathrm{obs}}$ has been required (where $N_{\mathrm{obs}}$ is the number of events in the population catalog)  [43, 46]. Then, if we discount the possibility of requiring more injections to be analyzed ad-hoc, in general at high computational cost, a large class of models cannot be studied using a generic (broadly/uniformly distributed) injection set [47]. For example, population models where $\pi(\theta|\Lambda)$ is strongly peaked around a specific point, if the density of injections $p_{\mathrm{draw}}(\theta)$ is featureless (e.g. [48]); or in general, where some $w_{i}$ are $\gg 1$ (see also [49]).
As an alternative, using an analytical function – or in general, a continuous function – that accurately describes the probability of detection of BBH signals will make it possible to estimate $P_{\mathrm{det}}$ without such restrictions on the model and original injection sampling distributions. As an initial effort towards such a description, the dependence on masses and spins of the selection function can be modelled via “semi-analytic” methods with various possible corrections. The interferometer network sensitivity is approximated by evaluating the expected (optimal) signal-to-noise ratios (SNRs) of signals for fixed, ideal detector power spectral densities (PSDs). The probability of detection is then obtained by assuming that the network is sensitive to CBCs with SNR above a fiducial threshold value [26, 50], which may be tuned to approximate the number of actual signals, or injections detected by search pipelines [51, 52].
However, a simple optimal SNR threshold is not necessarily an accurate approximation for $P_{\mathrm{det}}$, due to many factors which violate the assumptions behind the “semi-analytic” calculation. Beyond the effect of the noise realization on matched filter SNR (e.g. [52]), for instance, the number of detectors observing and their sensitivity varies over time [53, 54, 55, 56]; detector noise is also not completely described by a single set of PSDs, being not perfectly Gaussian and stationary. In reality there are excess noise transients (“glitches”, see e.g. [57, 58]) that may be partly suppressed by search pipelines employing additional quantities beyond the SNR, and other non-ideal noise behaviour [59, 60]. In general the search ranking statistic may have additional dependences beyond the SNR, for example explicit dependence on masses and/or spins [61, 62] [^2] Such dependence amounts to modelling a signal population for which the search is optimized [63]: like other aspects of the event ranking, this must be determined in advance of performing the search., and/or on detector outputs other than strain which are correlated with transient noise [64, 65, 66].
@@
The strategy we pursue here to overcome this limitation is to fit a continuous and physically motivated $P_{\mathrm{det}}$ function over distance, masses and, ultimately, spins. While the basic functional form is derived from the known theoretical dependence of SNR on distance and chirp mass, we introduce corrections to account for non-ideal detector behaviour, and for higher order effects of masses and spins on an event’s detectability in actual search pipelines. The fit is adjusted to maximize the likelihood of found injections in the O3 sensitivity data release corresponding to GWTC-3 [69, 3].
@@ 2 Methods
LIGO-Virgo-KAGRA (LVK) O3 BBH search sensitivity estimates [69] were used for the analysis. This data release contains sets of simulated signals (injections) from binary black holes that can be used to estimate the search sensitivity corresponding to the GWTC-3 catalog [70].
@@ 2.1 Sigmoid distance dependence model
Considering a source of fixed intrinsic parameters (binary component masses and spins), the dependence on distance of search sensitivity is mainly determined by the sky position and orientation of the source relative to a detector network (assumed to have stationary sensitivity), described by various angular parameters. If the angular parameters were also fixed, then the amplitude and expected SNR are simply inversely proportional to distance: then, assuming the source is detected if SNR is above a given threshold, $P_{\mathrm{det}}(d_{L})$ would be almost 1 for some range of distances up to a given point, and then close to 0 for larger distances, as represented in Fig. 1a. The detection probability has the shape of a sigmoid instead of a step function because of noise in the detector: in an actual search, detection will be determined by the matched filter SNR, a function of the data which contains a stochastic term resulting from the detector noise realization in addition to the expected SNR.
@@
Since the sources’ positions and orientations are a priori unknown, and isotropically distributed throughout the sky, we need to consider all the possible values of the angular parameters and average over them. As a result of this marginalization, the functional form of the search sensitivity takes a form more similar to Fig. 1b. These effects on search sensitivity were quantified by Finn & Chernoff (FC) [71], predicting the form of the probability of detection function over distance by considering geometrical factors in the expected SNR for the case of only one detector with fixed sensitivity.
Although the true functional dependence of $P_{\mathrm{det}}(d_{L})$ in a realistic case will have many corrections relative to this idealized calculation, our fundamental model assumption is that such corrections are small. Thus, as a starting point we use a parameterized sigmoid function that is able to fit the FC data ([71] Table I) faithfully, while also allowing for minor deviations. The shape of the sigmoid is controlled by four parameters $\varepsilon_{\max},\alpha,\delta$ and $\gamma$, out of which $\alpha=2.05$ is fixed to the value obtained in a direct fit to FC data and the other 3 will be free model parameters. The probability of detection, for a binary with given intrinsic parameters, thus reads as
@@
Having modelled the dependence on amplitude and distance for fixed intrinsic binary parameters, we then add a dependence on these intrinsic parameters by writing $d_{\mathrm{mid}}(m_{1z},m_{2z})$ and $\varepsilon_{\max}(m_{1z},m_{2z})$ as a function of the redshifted masses (we will also allow for component spin dependence in Sec. 4). In the low mass limit, $d_{\mathrm{mid}}$ is expected to take the form $d_{\mathrm{mid}}\sim D_{0}f_{d}\mathcal{M}_{z}^{5/6}$, where $D_{0}$ is a constant corresponding to average detector sensitivity over a given observing run; i.e. sensitive distance should scale with $\mathcal{M}_{z}^{5/6}$ for the simplest inspiral model, where the redshifted chirp mass is given by $\mathcal{M}_{z}=\eta^{3/5}M_{z}$ in terms of the redshifted binary total mass $M_{z}=m_{1z}+m_{2z}$ and the symmetric mass ratio $\eta=m_{1}m_{2}/M^{2}\equiv m_{1z}m_{2z}/M_{z}^{2}$. The remaining dependence on masses, corresponding to corrections to this lowest order inspiral SNR power law behaviour, is then to be fitted by a function $f_{d}(M_{z},\eta)$. Taking into account that $\varepsilon_{\max}$ should not be higher than 1 and always positive, we take a functional form $\varepsilon_{\max}\sim 1-\exp(f_{\varepsilon})$, with $f_{\varepsilon}$ a function of intrinsic parameters.
At this point, the functions providing mass dependence to $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ could be any arbitrary combination of $M_{z}$ and $\eta$. In order to gain more insight into the needed corrections, we perform mass binned fits to the detected injections released with the GWTC-3 catalog.
@@ 2.2 Mass binned fits
In order to gain insight into the functional forms of $f_{d}$ and $f_{\varepsilon}$ that will be able to fit injection results over the BBH mass space, we first fit our analytical model of $P_{\mathrm{det}}$ as a function of distance for fixed masses, or rather in mass bins, by numerically maximizing the likelihood of the injection campaign search results. [^3] Code to evaluate our $P_{\mathrm{det}}$ model and fit its parameters to injection results, and the resulting maximum likelihood parameter values, are publicly available at https://github.com/AnaLorenzoMedina/cbc_pdet. In this binned fit procedure, rather than being continuous functions, $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ are treated as scalar parameters, fitted separately for each mass bin; we choose bins equally spaced in the logarithms of $m_{1}$ and $m_{2}$, as they then contain comparable counts of found injections with FAR $\leq 1$/yr [14], at least for mass ratios close to unity. For a given mass bin labeled as $\cdot^{M}$ we use a inhomogeneous Poisson likelihood for the found injections:
@@
$\mu^{M}\equiv N_{\exp}=\int_{0}^{\infty}\lambda^{M}(d_{L})\,\mathrm{d}d_{L}\,.$ (7)
Although $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ vary between bins, we use the same shape parameter values $\delta$ and $\gamma$ over the whole mass range, since we assume $P_{\mathrm{det}}$ should have the same shape regardless of the mass. We then perform a fit in two alternating stages. In the ‘sigmoid fitting’ stage, $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ are held fixed for all bins and the sigmoid shape is fitted by maximizing the total log likelihood $\sum_{M}\ln\mathcal{L}^{M}$ over $\delta$ and $\gamma$. In the ‘mass fitting’ stage, $\delta$ and $\gamma$ are instead fixed and the values of $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ are fitted separately for each bin by maximizing $\ln\mathcal{L}^{M}$. The two stages are alternated until the total likelihood has converged to a global maximum (see e.g. [73] for a general discussion of “alternating optimization”). For optimization at each stage the Nelder-Mead simplex algorithm was used, a typical direct search method for multidimensional unconstrained minimization (e.g. [74]).
Figure 2: Optimized sensitive distance $d_{\mathrm{mid}}$ estimates from the mass binned fits. Points are colored according to the z-axis values $d_{\mathrm{mid}}/\mathcal{M}_{z}^{5/6}$ for clarity.
Figure 2 shows the resulting optimized values of $d_{\mathrm{mid}}$, having scaled out a factor of $\mathcal{M}_{z}^{5/6}$. While all the values are comparable in magnitude, we see strong linear and quadratic dependencies on $M_{z}$ at higher $\eta$, decreasing as the symmetric mass ratio gets smaller. Taking into account this behaviour, we first proposed the following expression to describe the global mass dependence: [^4] The $a$ coefficient subscripts, in order, represent powers of $M_{z}$ and $(1-4\eta)$.
$\;\;d_{\mathrm{mid}}^{*}(m_{1z},m_{2z})=D_{0}\mathcal{M}_{z}^{5/6}\left[1+a_{10}M_{z}+a_{20}M_{z}^{2}+(a_{01}+a_{11}M_{z}+a_{21}M_{z}^{2})(1-4\eta)\right],$ (8)
in which we are separating out the dependence away from equal masses (i.e. for $\eta<1/4$) where there may be higher uncertainty, treating this as a correction for unequal mass binaries. This first fit was already used in a population analysis to investigate the evolution of the BBH population over redshift [21]. However this function has the disadvantage that it may become negative, which would be physically disallowed; therefore for our main results we use a slightly modified form with corrections via an exponential:
@@
We next consider the functional dependence of $\varepsilon_{\max}$ (the detection probability in the limit of zero distance) by inspecting the optimized values of $\varepsilon_{\max}$ for binned fits in Fig. 3. Here, we allow (in principle unphysical) $\varepsilon_{\max}$ values greater than 1 which may be caused by small number statistics; although the total number of found injections in a bin may be large, due to their distribution over distance, the detection probability at small distance is typically badly constrained. The trend of variation is most apparent over linear total mass, therefore given the restrictions stated previously we chose the following function:
@@
Figure 3: Maximum search efficiency $\varepsilon_{\max}$ fitted values for each mass bin: the color bar shows the number of detected events (injections) in each bin.